Using Gauss’ theorem calculate the flux of the vector field A = x³i + y³I + z³k
through the surface of a hemisphere of radius A which has the centre of its base at the
origin.
1
Expert's answer
2020-06-08T10:29:06-0400

The Gauss' theorem is "\iint_S(A\sdot n)ds=\iiint_V(\nabla \sdot A)dv" as "\nabla\sdot A=3(x^2+y^2+z^2)" then

"\iiint_V(\nabla\sdot A)dv=3\sdot4\int_{0}^{\pi/2}d\phi\int_{0}^{\pi/2}\sin\theta d\theta\int_{0}^{A}r^2dr="

"=12\pi/2(-\cos\theta)|_0^{\pi/2}A^3/3=2\pi A^3"